tan 100° + tan 125° + tan 100°tan125°

Grade:11

2 Answers

Vijay Mukati
8 years ago
Hint: Please use the formula for tan(A+B) to simplify it. Also please note that tan(100+125) = tan(225) = tan(180 +45) = tan45 =1.
Aakash
8 years ago
\tan (100+125) = \frac{\tan (100) + \tan (125))}{1-\tan (100)\tan (125)}
 
\tan (225) = \frac{\tan (100) + \tan (125))}{1-\tan (100)\tan (125)}
 
=\tan (180+45) = \frac{\tan (100) + \tan (125))}{1-\tan (100)\tan (125)}
 
=\tan (45) = \frac{\tan (100) + \tan (125))}{1-\tan (100)\tan (125)}
 
1 = \frac{\tan (100) + \tan (125))}{1-\tan (100)\tan (125)}
 
1-\tan (100)\tan (125) =\tan (100) +\tan (125)
 
\tan (100) +\tan (125)-\tan (100)\tan (125) = 1
 
PLEASE APPROVE